# Holographic entanglement negativity and thermodynamics in backreacted AdS black hole

### v1

@@ Abstract
We investigate holographic entanglement negativity (HEN) as a probe of mixed state quantum correlations in a deformed AdS black hole background with backreaction sourced by a string cloud. The bulk geometry is dual to a strongly coupled large-$N_{c}$ gauge theory at finite temperature, backreacted by a finite density of heavy static fundamental quarks. Alongside holographic entanglement entropy (HEE), we analyze entanglement thermodynamics and establish a first law like relation in the small backreaction regime by identifying the deformed black hole at zero temperature as the natural ground state of the dual field theory. Using analytic expansions in the low and high temperature limits, we compute the HEN for adjacent, disjoint and bipartite subsystem configurations and examine its dependence on the backreaction parameter $\rho$. Our results reveal a clear regime dependence: at low temperature, backreaction suppresses the negativity, signaling a reduction of distillable quantum correlations while at high temperature the negativity is enhanced due to the monotonic $\rho$-dependence of near horizon contributions. We further compare the behavior of HEN with HEE and mutual information, demonstrating that entanglement negativity provides a sharper diagnostic of the competition between thermal and quantum correlations in holographic plasmas with matter backreaction.
@@ 1 Introduction
Strongly coupled quantum systems appear in a wide range of physical settings ranging from condensed matter systems and cold atomic gases to the quark–gluon plasma formed in heavy-ion collisions. A central difficulty in analyzing such systems is the characterization of their quantum correlations, particularly in regimes where perturbative methods fail. The AdS/CFT correspondence [Maldacena:1997re, Witten:1998qj, Gubser:1998bc] offers a powerful non-perturbative framework by relating strongly coupled gauge theories to weakly coupled gravitational theories in higher dimensional spacetimes. Within this holographic framework, quantum information theoretic measures such as entanglement entropy, mutual information, and computational complexity have emerged as valuable probes of the correlation structure of the dual field theories. Among these, the holographic computation of entanglement entropy through the Ryu–Takayanagi (RT) prescription and its covariant generalization [Ryu:2006bv, Hubeny:2007xt] has played a central role. The RT prescription has found extensive application in investigating the entanglement entropy of conformal field theories at finite temperature which are holographically dual to AdS-Schwarzschild black holes [Cadoni:2010ztg, Solodukhin:2006xv, Hubeny:2012ry, Fischler:2012ca, Fischler:2012uv]. Detailed analyses of the resulting low and high temperature behaviors of holographic entanglement entropy led to the formulation of a thermodynamics like first law for the entanglement entropy in [Bhattacharya:2012mi, Allahbakhshi:2013rda]. Entanglement thermodynamics has also been investigated for charged black holes in AdS4 backgrounds. In particular, the effects of temperature and charge on the entanglement entropy of the boundary field theory dual to charged AdS black holes were studied in [Chaturvedi:2016kbk], with extensions to higher-dimensional charged black holes presented in [Karar:2018ecr]. Furthermore, entanglement thermodynamics has been explored in a variety of other scenarios including non-conformal and non-relativistic geometries [Blanco:2013joa, Dey:2014voa, Mansoori:2015sit, Park:2015afa, Chakraborty:2014lfa, Mishra:2015cpa, McCarthy:2017amh, OBannon:2016exv, Sun:2016til, Nadi:2019bqu, Saha:2019ado, Maulik:2020tzm].
@@
The entanglement entropy does not provide a faithful characterization of quantum correlations in mixed states as it captures both quantum and classical contributions. To extract the purely quantum correlations in such states, entanglement negativity was introduced in quantum information theory [Vidal:2002zz, Plenio:2005cwa]. Entanglement negativity serves as computable upper bound on distillable entanglement, and monotonic under local operations and classical communication (LOCC). Further it vanishes for separable states, thereby making it a robust measure of quantum correlations in mixed state scenarios. In recent years, entanglement negativity has gained prominence as a powerful diagnostic of quantum correlations in quantum field theories, capturing information inaccessible through entanglement entropy alone. A field-theoretic formulation based on the replica trick was developed in [Calabrese:2012nk] enabling the computation of negativity in extended systems. This approach was later applied to finite temperature conformal field theories [Calabrese:2014yza] and to lattice models such as the critical Ising chain [Calabrese:2013mi]. The time-dependent behavior of entanglement negativity have also been explored in both local and global quantum quenches revealing universal features of entanglement spreading in conformal field theories [Wen:2015qwa, Coser:2014gsa]. These results collectively establish entanglement negativity as a central tool in characterizing quantum correlations in many body systems. Motivated by these developments, several holographic prescriptions for entanglement negativity have been proposed [Rangamani:2014ywa, Chaturvedi:2016rcn, Chaturvedi:2017znc, Jain:2017aqk, Jain:2017uhe, Malvimat:2018txq, Malvimat:2018ood, Chaturvedi:2016rft, Jain:2017xsu, Jain:2018bai, KumarBasak:2020viv, Mondal:2021kzj, Afrasiar:2021hld, Kudler-Flam:2018qjo, Kusuki:2019zsp, KumarBasak:2020eia, KumarBasak:2021lwm] extending the intuition gained from the RT prescription. The HEN turns out to be an effective measure to study the behavior of entanglement in the vicinity of critical points for strongly coupled gauge theory [Karan:2023hfk]. These studies reveal that negativity is sensitive to critical behavior and captures features that are not visible through entanglement entropy alone. In parallel, HEN provides a robust characterization of quantum correlations in the presence of thermal and deformation effects in quantum field theories at finite temperature [Pant:2024eno]. Together, these developments underscore the effectiveness of HEN as a better probe of mixed state entanglement in holographic frameworks. Closely related insights arise from the study of reflected entropy, that offers an alternative mixed state measure based on the entanglement structure of canonical purifications [Dutta:2019gen]. While reflected entropy captures both quantum and classical correlations, its holographic dual via the entanglement wedge cross section provides complementary geometric information that enriches the understanding of correlation structure beyond what is accessible through negativity alone. Together, holographic entanglement negativity and reflected entropy describes a broad framework for probing mixed state entanglement in strongly coupled dual field theories.
@@
Recent studies have highlighted the significance of investigating entanglement measures in bulk geometries that include matter backreaction as such effects capture additional physical features of the dual strongly coupled field theories. An instructive example is provided by a deformed AdS black hole geometry sourced by a homogeneous string cloud [Chakrabortty:2011sp] which is holographically dual to a large-$N_{c}$ gauge theory at finite temperature backreacted by a uniform density of static heavy quarks. The backreaction parameter encodes the quark density and modifies both the near boundary and near horizon geometry, providing a controlled setup to investigate how heavy-flavor matter influences entanglement structure, thermodynamics, and mixed-state quantum correlations. This setup has been extensively employed to explore transport properties, entanglement structure and chaotic behavior in the dual field theory [Chakrabortty:2016xcb, Chakrabortty:2020ptb, Chakrabortty:2022kvq, Jain:2023xta].
@@
A natural question that arises is whether holographic entanglement entropy admits a relation similar to the first law of thermodynamics in such backreacted field theories. To address this, we study entanglement thermodynamics in deformed strongly coupled field theories in the context of AdS4/CFT3. Specifically, we analyze the dependence of the entanglement entropy on temperature and backreaction for a strip shaped subsystem in the boundary field theory dual to a deformed AdS4 black hole. Furthermore, while the impact of backreaction on entanglement entropy and related probes has been widely investigated, the behavior of HEN in this background remains largely unexplored. Given that entanglement negativity provides a better probe of distillable quantum entanglement in mixed states, a systematic analysis in this context is both natural and timely. In this work, we present a detailed analysis of HEN for adjacent, bipartite and disjoint subsytems in the the string cloud deformed AdS black hole geometry. Employing analytic expansions in both low and high temperature regimes, we find a clear dichotomy in the effect of backreaction. At low temperature, the leading $\rho$-dependent correction enters as $(1-\rho)T_{f}^{\,d}$ and suppresses the HEN which indicates the inclusion of heavy static quarks enhances thermal correlations at the cost of quantum correlations. By contrast, in the high temperature regime, the HEN scales with $T_{f}$ and is enhanced with increasing $\rho$, as the near horizon contributions encoded in the $\widetilde{S}(\rho,d)$ coefficients grow monotonically with the quark density. This dichotomy persists for adjacent and disjoint intervals while in the bipartite configuration, the interplay between $\rho T_{f}^{\,d-1}$ and $(1-\rho)T_{f}^{d}$ terms leads to richer, parameter-dependent behaviour. A comparison with HEE shows that while, HEE is dominated by thermal scaling and is insensitive to the separation of quantum and classical correlations, HEN captures the genuine quantum contribution and thus provides a sharper diagnostic of backreaction effects in strongly coupled plasmas. Further, HEN for two disjoint subsystems allow us to compute the critical separation at which the distillable entanglement vanishes. We give a master equation that can be used to find the critical separation for particular values of backreaction and dimensions.
@@
The structure of the paper is as follows. In section˜2, we introduce the string cloud deformed AdS black hole geometry and its holographic dual interpretation. Section 3 provides a review of holographic entanglement entropy in this background, with particular emphasis on the analytic behavior in the low and high temperature regimes. In section˜4, we analyze the holographic entanglement entropy in boundary field theories dual to deformed AdS4 black holes at both zero and finite temperatures in the regime of small backreaction, and and establish a first law of entanglement thermodynamics at low temperatures. In section˜5, we present the computation of holographic entanglement negativity for adjacent subsystems followed by the analysis of bipartite configuration in section˜6. The HEN for the configuration of two disjoint subsystems is examined in section˜7 followed by a critical separation analysis in section˜8. Finally, in section˜9, we summarize our findings.
2 Background
@@
We begin by briefly reviewing the holographic setup corresponding to a strongly coupled gauge theory with a finite density of heavy static fundamental quarks, which is dual to a deformed AdS black hole geometry. This deformation arises due to the backreaction of a homogeneous distribution of long, non-interacting, and static strings stretched along the radial direction in the bulk spacetime. The endpoints of these strings are attached to the AdS boundary and represent static fundamental quarks, while the string bodies encode the associated gluonic degrees of freedom [Chakrabortty:2011sp].
@@
As suggested in [Chakrabortty:2011sp], the thermodynamic limit ($V_{d-1}\to\infty$), to keeps $N/V_{d-1}$ finite for maintaining constant quark density as $N\to\infty$. The only non-vanishing components of the energy-momentum tensor are,
@@
$C=\frac{\partial M}{\partial T}=\frac{V_{d-1}(d-1)R^{2d-2}(d(d-1)R^{2d}z_{H}-2bR^{4}z_{H}^{d})}{4G_{d+1}z_{H}^{d-1}(d(d-1)R^{2d}z_{H}+2(d-2)bR^{4}z_{H}^{d})}$ (2.15)
@@
${\cal F}=-\frac{(d-1)z_{H}+2bR^{2}(d-2)z_{H}^{d}}{16\pi z_{H}^{d+1}(d-1)}.$ (2.18)
To examine the dynamical (geometric) stability of the background spacetime, one can carry out a gauge-invariant analysis of gravitational perturbations using the formalism developed in [Ishibashi:2011ws, Kodama:2003jz, Ishibashi:2003ap, Kodama:2003kk, Kodama:2007ph]. Within this approach, the background spacetime is regarded as a product manifold:
@@
Here, ${\mathcal{N}}^{2}$ denotes a two-dimensional spacetime with coordinates $(t,z)$, while ${\mathcal{K}}^{p}$ represents a $p$-dimensional maximally symmetric space spanned by the spatial coordinates $x^{i}$. Gravitational perturbations are classified into scalar, vector, and tensor modes on ${\mathcal{K}}^{p}$, which allows the Einstein equations to be analyzed in a decoupled manner. The tensor and vector sectors have been studied explicitly in [Chakrabortty:2011sp], where the background geometry was shown to be stable against these classes of perturbations.
@@ 3 Revisiting the holographic entanglement entropy and EWCS
Since entanglement thermodynamics and holographic entanglement negativity are primarily studied in this work, it is natural to first review the computation of holographic entanglement entropy in the backreacted AdS black hole geometry, as introduced in section˜2. Moreover, as reflected entropy is directly related to the entanglement wedge cross section, we also briefly review the EWCS. Both quantities were originally computed in [Chakrabortty:2020ptb].
3.1 Holographic entanglement entropy (HEE)
@@
$S_{A}=\frac{L^{d-2}R^{d-1}}{2G^{d+1}_{N}}\int_{0}^{z_{t}}\frac{dz}{z^{d-1}\sqrt{h(z)}\sqrt{1-\left(\frac{z}{z_{t}}\right)^{2d-2}}},$ (3.7)
To obtain the HEE in terms of boundary parameters, it is necessary to express the turning point $z_{t}$ as a function of the strip width and temperature. However, evaluating the integral in eq.˜3.6 for arbitrary dimension $d$ and finite temperature is analytically intractable. As discussed in [Chakrabortty:2020ptb], we therefore analyze the problem in two regimes i.e. the low and high temperature limits where the integrals can be computed explicitly, and subsequently use these results to extract the corresponding behavior of the HEE.
Low-temperature expansion:
By taking the low temperature limit i.e $T_{f}l\ll 1$ ( or $z_{t}/z_{H}\ll 1$) the expression of $z_{t}$ becomes
@@
Here $S_{0},S_{1}$ and $S_{2}$ are terms that only depends on dimensions parameter $d$, and are given by
@@
Finally the HEE in low temperature limit is given by
$S_{A}^{\mathrm{low}}=\frac{R^{d-1}L^{d-2}}{2G_{N}}\left[\frac{1}{(d-2)\epsilon^{d-2}}+S_{0}\,l^{2-d}+S_{0}S_{2}\,\rho\,l\left(\frac{4\pi T_{f}}{d}\right)^{d-1}+S_{0}S_{1}(1-\rho)\,l^{2}\left(\frac{4\pi T_{f}}{d}\right)^{d}+\cdots\right].$ (3.11)
The finite pieces $S_{0}\,l^{2-d}$, $S_{0}S_{2}\,\rho\,lT_{f}^{d-1}$, and $S_{0}S_{1}(1-\rho)l^{2}T_{f}^{d}$, are exactly the contributions that will appear in the HEN formulas.
@@
In eq.˜3.11, the first term is divergent containing UV-cut off ($\epsilon$). In the low-temperature regime, the entanglement entropy exhibits nontrivial modifications due to backreaction. Unlike the charged plasma case [Kundu:2016dyk], where the leading correction scales as $T_{f}^{\,d}$, here the dominant term is proportional to $T_{f}^{\,d-1}$. This distinction arises from the blackening function $h(z)$: near the boundary the $z^{\,d-1}$ contribution from the quark density dominates over the mass term $\sim z^{d}$, making backreaction the leading source of corrections to the HEE at low effective temperature. It will be particularly interesting to examine whether a similar scaling behavior also manifests in the computation of HEN.
High-temperature expansion:
@@
In the high temperature regime $T_{f}l\gg 1$ (or $z_{t}\to z_{H}$), the RT surface approaches the horizon. Although, the direct series expansions for $l$ and $A$ diverge in this limit, an appropriate combination of these quantities yields a finite and well behaved expression for the entanglement entropy as
$\mathcal{A}-\frac{L^{d-2}R^{d-1}}{z_{t}^{\,d-1}}\,l=\frac{2L^{d-2}R^{d-1}}{z_{t}^{\,d-2}}\int_{0}^{1}\frac{du}{\sqrt{h(u)}\,u^{\,d-1}\sqrt{1-u^{2d-2}}}-\cdots,$ (3.12)
where $u=\frac{z}{z_{t}}$. This leads to the finite part of HEE as
$S^{high}_{A}\simeq\frac{R^{d-1}}{4G_{d+1}}\left[V\left(\frac{4\pi T_{f}}{d}\right)^{\!d-1}\left(1+2\frac{d}{4\pi T_{f}l}\,\widetilde{S}(\rho,d)\right)\right],$ (3.13)
where $V=lL^{d-2}$ and
$\widetilde{S}(\rho,d)=\frac{\sqrt{\pi}\,\Gamma\left(-\frac{d-2}{2d-2}\right)}{(2d-2)\,\Gamma\left(\frac{1}{2d-2}\right)}+\int_{0}^{1}du\left[\frac{\sqrt{1-u^{2d-2}}}{u^{\,d-1}\sqrt{h(u)}}-\frac{1}{u^{\,d-1}\sqrt{1-u^{2d-2}}}\right].$ (3.14)
@@
At high $T_{f}$, the leading term reproduces the thermal entropy density via its volume scaling, with $\rho$ entering only at subleading order.
@@
Across both the low and high effective temperature limits, the quark density parameter $\rho$ plays a crucial role in determining the structure of the finite part of the entanglement entropy. At low $T_{f}$, the dominant correction scales as $T_{f}^{\,d-1}$, which originates from the backreaction-induced $z^{\,d-1}$ contribution through $h(z)$. In contrast, at high $T_{f}$ the leading term coincides with the thermal entropy density while the dependence on $\rho$ manifests only through subleading corrections.

### v2

@@ Abstract
We investigate holographic entanglement negativity (HEN) as a probe of mixed state quantum correlations in a deformed AdS black hole background with backreaction sourced by a string cloud. The bulk geometry is dual to a strongly coupled large-$N_{c}$ gauge theory at finite temperature, backreacted by a finite density of heavy static fundamental quarks. We analyze entanglement thermodynamics and establish a first law like relation in the small backreaction regime by identifying the deformed black hole at zero temperature as the natural ground state of the dual field theory. Using analytic expansions in the low and high temperature limits, we compute the HEN for adjacent, bipartite and disjoint subsystem configurations and examine its dependence on the backreaction parameter $\rho$. Our results reveal that at low temperature and high effective temperature, backreaction increases the HEN signaling enhancement of distillable quantum correlations. We further compare the behavior of HEN with holographic entanglement entropy (HEE) and mutual information (MI), demonstrating that entanglement negativity provides a sharper diagnostic of the competition between thermal and quantum correlations in holographic plasmas with matter backreaction.
@@ 1 Introduction
Strongly coupled quantum systems appear in a wide range of physical settings ranging from condensed matter systems and cold atomic gases to the quark–gluon plasma formed in heavy-ion collisions. A central difficulty in analyzing such systems is the characterization of their quantum correlations, particularly in regimes where perturbative methods fail. The AdS/CFT correspondence [58, 83, 29] offers a powerful non-perturbative framework by relating strongly coupled gauge theories to weakly coupled gravitational theories in higher dimensional spacetimes. Within this holographic framework, quantum information theoretic measures such as entanglement entropy, mutual information, and computational complexity have emerged as valuable probes of the correlation structure of the dual field theories. Among these, the holographic computation of entanglement entropy through the Ryu–Takayanagi (RT) prescription and its covariant generalization [73, 30] has played a central role. The RT prescription has found extensive application in investigating the entanglement entropy of conformal field theories at finite temperature which are holographically dual to AdS-Schwarzschild black holes [8, 76, 31, 28, 27]. Detailed analyses of the resulting low and high temperature behaviors of holographic entanglement entropy led to the formulation of a thermodynamics like first law for the entanglement entropy in [6, 2]. Entanglement thermodynamics has also been investigated for charged black holes in AdS4 backgrounds. In particular, the effects of temperature and charge on the entanglement entropy of the boundary field theory dual to charged AdS black holes were studied in [17], with extensions to higher-dimensional charged black holes presented in [46]. Furthermore, entanglement thermodynamics has been explored in a variety of other scenarios including non-conformal and non-relativistic geometries [7, 23, 61, 70, 16, 64, 55, 63, 68, 77, 66, 74, 62].
@@
The entanglement entropy does not provide a faithful characterization of quantum correlations in mixed states as it captures both quantum and classical contributions. To extract the purely quantum correlations in such states, entanglement negativity was introduced in quantum information theory [81, 71]. Entanglement negativity serves as computable upper bound on distillable entanglement, and monotonic under local operations and classical communication (LOCC). Further it vanishes for separable states, thereby making it a robust measure of quantum correlations in mixed state scenarios. In recent years, entanglement negativity has gained prominence as a powerful diagnostic of quantum correlations in quantum field theories, capturing information inaccessible through entanglement entropy alone. A field-theoretic formulation based on the replica trick was developed in [9] enabling the computation of negativity in extended systems. This approach was later applied to finite temperature conformal field theories [10] and to lattice models such as the critical Ising chain [11]. The time-dependent behavior of entanglement negativity have also been explored in both local and global quantum quenches revealing universal features of entanglement spreading in conformal field theories [82, 21]. These results collectively establish entanglement negativity as a central tool in characterizing quantum correlations in many body systems. Motivated by these developments, several holographic prescriptions for entanglement negativity have been proposed [72, 20, 18, 40, 39, 60, 59, 19, 37, 38, 54, 65, 1, 50, 56, 52, 53] extending the intuition gained from the RT prescription. The holographic proposals for entanglement negativity based on algebraic sums of bulk geodesics were further substantiated in [51] for subsystems with spherical entangling surfaces through the analysis of replica symmetry breaking saddles in the gravitational path integral [25]. More recently, [24] has introduced a generalized framework for the holographic dual of negativity for general holographic states based on the modified cosmic brane proposal. The HEN turns out to be an effective measure to study the behavior of entanglement in the vicinity of critical points for strongly coupled gauge theory [45]. These studies reveal that negativity is sensitive to critical behavior and captures features that are not visible through entanglement entropy alone. In parallel, HEN provides a robust characterization of quantum correlations in the presence of thermal and deformation effects in quantum field theories at finite temperature [69]. Together, these developments underscore the effectiveness of HEN as a better probe of mixed state entanglement in holographic frameworks. Closely related insights arise from the study of reflected entropy, that offers an alternative mixed state measure based on the entanglement structure of canonical purifications [26]. While reflected entropy captures both quantum and classical correlations, its holographic dual via the entanglement wedge cross section provides complementary geometric information that enriches the understanding of correlation structure beyond what is accessible through negativity alone. Together, holographic entanglement negativity and reflected entropy describes a broad framework for probing mixed state entanglement in strongly coupled dual field theories.
@@
Recent studies have highlighted the significance of investigating entanglement measures in bulk geometries that include matter backreaction as such effects capture additional physical features of the dual strongly coupled field theories. An instructive example is provided by a deformed AdS black hole geometry sourced by a homogeneous string cloud [15] which is holographically dual to a large-$N_{c}$ gauge theory at finite temperature backreacted by a uniform density of static heavy quarks. The backreaction parameter encodes the quark density and modifies both the near boundary and near horizon geometry, providing a controlled setup to investigate how heavy-flavor matter influences entanglement structure, thermodynamics and mixed-state quantum correlations. This setup has been extensively employed to explore transport properties, entanglement structure and chaotic behavior in the dual field theory [12, 14, 13, 41].
@@
A natural question that arises is whether holographic entanglement entropy admits a relation similar to the first law of thermodynamics in such backreacted field theories. To address this, we study entanglement thermodynamics in deformed strongly coupled field theories in the context of AdS4/CFT3. Specifically, we analyze the dependence of the entanglement entropy on temperature and backreaction for a strip shaped subsystem in the boundary field theory dual to a deformed AdS4 black hole. Furthermore, while the impact of backreaction on entanglement entropy and related probes has been widely investigated, the behavior of HEN in this background remains unexplored. Given that entanglement negativity provides a better probe of distillable quantum entanglement in mixed states, a systematic analysis in this context is both natural and timely. Therefore, in this work we present a detailed analysis of HEN for adjacent, bipartite and disjoint configurations in strongly coupled field theories dual to the string cloud deformed AdS black hole geometry. In our analysis we find that at low and high effective temperature regime, the HEN increases with backreaction in all three configurations at leading order. This observation supports the findings that extra degrees of freedoms introduced due to matter enhances the correlation in the strongly coupled field theory. Further, HEN for two disjoint subsystems allow us to compute the critical separation at which the distillable entanglement vanishes. In the leading order we observe that increasing backreaction increases the critical separation at which the negativity vanishes.
@@
The structure of the paper is as follows. In section˜2, we introduce the string cloud deformed AdS black hole geometry and its holographic dual interpretation. In section˜3 we provide a review of holographic entanglement entropy in this background, with particular emphasis on the analytic behavior in the low and high temperature regimes. In section˜4, we analyze the holographic entanglement entropy in boundary field theories dual to deformed AdS4 black holes at both zero and finite temperatures in the regime of small backreaction, and and establish a first law of entanglement thermodynamics at low temperatures. In section˜5, we present the computation of holographic entanglement negativity for adjacent subsystems followed by the analysis of bipartite configuration in section˜6. The HEN for the configuration of two disjoint subsystems is examined in section˜7 followed by a critical separation analysis in section˜8. Finally, in section˜9, we summarize our findings.
2 Deformed/Backreacted AdS Black Hole Spacetime
@@
We begin by briefly reviewing the holographic setup corresponding to a strongly coupled gauge theory with a finite density of heavy static fundamental quarks, which is dual to a deformed AdS black hole geometry. This deformation arises due to the backreaction of a homogeneous distribution of long, non-interacting, and static strings stretched along the radial direction in the bulk spacetime. The endpoints of these strings are attached to the AdS boundary and represent static fundamental quarks, while the string bodies encode the associated gluonic degrees of freedom [15].
@@
As suggested in [15], the thermodynamic limit ($V_{d-1}\to\infty$), to keeps $N/V_{d-1}$ finite for maintaining constant quark density as $N\to\infty$. The only non-vanishing components of the energy-momentum tensor are,
@@
$C=\frac{\partial M}{\partial T}=\frac{V_{d-1}(d-1)R^{2d-2}(d(d-1)R^{2d}z_{H}-2bR^{4}z_{H}^{d})}{4G_{N}^{d+1}z_{H}^{d-1}(d(d-1)R^{2d}z_{H}+2(d-2)bR^{4}z_{H}^{d})}$ (2.15)
@@
${\cal F}=-\frac{(d-1)R^{2d}z_{H}+2bR^{4}(d-2)z_{H}^{d}}{16\pi z_{H}^{d+1}(d-1)}.$ (2.18)
To examine the dynamical (geometric) stability of the background spacetime, one can carry out a gauge-invariant analysis of gravitational perturbations using the formalism developed in [33, 47, 32, 48, 49, 57]. Within this approach, the background spacetime is regarded as a product manifold:
@@
Here, ${\mathcal{N}}^{2}$ denotes a two-dimensional spacetime with coordinates $(t,z)$, while ${\mathcal{K}}^{p}$ represents a $p$-dimensional maximally symmetric space spanned by the spatial coordinates $x^{i}$. Gravitational perturbations are classified into scalar, vector, and tensor modes on ${\mathcal{K}}^{p}$, which allows the Einstein equations to be analyzed in a decoupled manner. The tensor and vector sectors have been studied explicitly in [15], where the background geometry was shown to be stable against these classes of perturbations.
@@ 3 Revisiting the Holographic Entanglement Entropy and EWCS
Since entanglement thermodynamics and holographic entanglement negativity are primarily studied in this work, it is natural to first review the computation of holographic entanglement entropy in the backreacted AdS black hole geometry, as introduced in section˜2. Moreover, as reflected entropy is directly related to the entanglement wedge cross section, we also briefly review the EWCS. Both quantities were originally computed in [14].
3.1 Holographic Entanglement Entropy
@@
$S_{A}=\frac{L^{d-2}R^{d-1}}{2G^{d+1}_{N}}\int_{0}^{z_{t}}\frac{dz}{z^{d-1}\sqrt{h(z)}\sqrt{1-\left(\frac{z}{z_{t}}\right)^{2d-2}}}.$ (3.7)
To obtain the HEE in terms of boundary parameters, it is necessary to express the turning point $z_{t}$ as a function of the strip width and temperature. However, evaluating the integral in eq.˜3.6 for arbitrary dimension $d$ and finite temperature is analytically intractable. As discussed in [14], we therefore analyze the problem in two regimes i.e. the low and high effective temperature limits where the integrals can be computed explicitly, and subsequently use these results to extract the corresponding behavior of the HEE.
Low temperature expansion:
By taking the low effective temperature limit i.e $T_{f}l\ll 1$ ( or $z_{t}/z_{H}\ll 1$) the expression of $z_{t}$ becomes
@@
Here $S_{0},S_{1}$ and $S_{2}$ are only function of dimension parameter $d$, and are given by
@@
These coefficients are negative for particular values of $d$. Finally, the HEE in low effective temperature limit is given by [14]
$S_{A}^{\mathrm{low}}=\frac{R^{d-1}L^{d-2}}{4G_{N}^{d+1}}\left[\frac{2}{(d-2)\epsilon^{d-2}}+S_{0}\,l^{2-d}+S_{0}S_{2}\,\rho\,l\left(\frac{4\pi T_{f}}{d}\right)^{d-1}+S_{0}S_{1}(1-\rho)\,l^{2}\left(\frac{4\pi T_{f}}{d}\right)^{d}+\cdots\right]$ (3.11)
The finite pieces $S_{0}\,l^{2-d}$ is negative contribution while $S_{0}S_{2}\,\rho\,lT_{f}^{d-1}$ is positive. On the other hand $S_{0}S_{1}(1-\rho)l^{2}T_{f}^{d}$ depends on the value of $\rho$. Therefore, the leading terms involving backreaction has positive contribution while the subleading term has reducing effect with respect to backreaction. Since we are in low temperature regime the dominating term with backreaction is $T_{f}^{d-1}$ that suggests enhancement of HEE with respect to backreaction.
@@
In eq.˜3.11, the first term is divergent containing UV-cut off ($\epsilon$). In the low-temperature regime, the entanglement entropy exhibits nontrivial modifications due to backreaction. Unlike the charged plasma case [55], where the leading correction scales as $T_{f}^{\,d}$, here the dominant term is proportional to $T_{f}^{\,d-1}$. This distinction arises from the blackening function $h(z)$ near the boundary the $z^{\,d-1}$ contribution from the string density dominates over the mass term $\sim z^{d}$, making backreaction the leading source of corrections to the HEE at low effective temperature. It is also observed that this term follows a volume law scaling which suggests that the backreaction also introduces extensive classical correlations (thermal-like) into the subsystem.
High temperature expansion:
@@
In the high effective temperature regime $T_{f}l\gg 1$ (or $z_{t}\to z_{H}$), the RT surface approaches the horizon. Although, the direct series expansions for $l$ and $A$ diverge in this limit, an appropriate combination of these quantities yields a finite and well behaved expression for the entanglement entropy as
$\mathcal{A}-\frac{L^{d-2}R^{d-1}}{z_{t}^{\,d-1}}\,l=\frac{2L^{d-2}R^{d-1}}{z_{t}^{\,d-2}}\int_{0}^{1}\frac{\sqrt{h(u)}\,(1-u^{2d-2})du}{u^{\,d-1}\sqrt{1-u^{2d-2}}},$ (3.12)
where $u=\frac{z}{z_{t}}$. This leads to the HEE at high temperature limit as [14]
$S^{high}_{A}=\frac{R^{d-1}}{4G_{N}^{d+1}}\left[\frac{2}{(d-2)}\left(\frac{L}{\epsilon}\right)^{d-2}+V\left(\frac{4\pi T_{f}}{d}\right)^{\!d-1}\left(1+2\left(\frac{d}{4\pi T_{f}l}\right)\,\widetilde{S}(\rho,d)\right)\right],$ (3.13)
where $V=lL^{d-2}$ is the volume of the strip subsystem $A$ in dual field theory and
$\widetilde{S}(\rho,d)=\frac{\sqrt{\pi}\,\Gamma\left(-\frac{d-2}{2d-2}\right)}{(2d-2)\,\Gamma\left(\frac{1}{2d-2}\right)}+\int_{0}^{1}du\left[\frac{\sqrt{1-u^{2d-2}}}{u^{\,d-1}\sqrt{h(z_{H}u)}}-\frac{1}{u^{\,d-1}\sqrt{1-u^{2d-2}}}\right].$ (3.14)
@@
At high $T_{f}$, the leading term reproduces the thermal entropy density via its volume scaling with $\rho$ entering only at subleading order. Note that the quantity $\widetilde{S}(\rho,d)$ is independent of $l$ in the high temperature limit as $z_{t}\to z_{H}$ and depends only on the backreaction parameter $\rho$ and the dimension $d$.
@@
Across both the low and high effective temperature limits, the quark density parameter $\rho$ plays a crucial role in determining the structure of the finite part of the entanglement entropy. At low $T_{f}$, the dominant correction scales as $T_{f}^{\,d-1}$ and volume of the subsystem which originates from the backreaction-induced $z^{\,d-1}$ contribution through $h(z)$. In contrast, at high $T_{f}$ the leading term coincides with the thermal entropy density while the dependence on $\rho$ manifests only through subleading corrections.